You want to synthetically create a one-year 100,000 tenge loan for yourself two years from now. You note that banks will lend today at the rate of 14% for two years and 12% for three years. Banks also will take time-deposits at 9% for two years and 10% for three years.
Answer #1

Inflow 2 years from now 100000

Outflow 3 years from now

Deposit now for 2 years at 9% 100000/1.09^2=84167.999327
Borrow now 84167.999327 for 3 years at 12%
Amount to be returned 3 years later=84167.999327*1.12^3=118249.978958